OpenStudy (anonymous):

Use the substitution method to solve the system of equations. 2x - 7y = 1 x - 8y = -4

OpenStudy (anonymous):

a) (2, 4) b) (4, 2) c) (1, 4) d) (4, 1)

OpenStudy (luigi0210):

solve for x first

OpenStudy (anonymous):

how?

OpenStudy (luigi0210):

Get it by itself

OpenStudy (anonymous):

i dont know how to solve for x?

OpenStudy (luigi0210):

well we are in trouble aren't we?

OpenStudy (campbell_st):

rewrite the 2nd equation with x as the subject then substitute that into equation 1 for x

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

2(1)-7Y=1 1-8y=-4

OpenStudy (anonymous):

are those the equations?

OpenStudy (whpalmer4):

doing as @campbell_st suggests: \[x-8y=-4\]\[x-8y+8y=-4+8y\]\[x=-4+8y\] Now rewrite the first equation, substituting (-4+8y) wherever you encounter \(x\). \[2x-7y=1\]\[2(-4+8y)-7y=1\]Distribute the 2 \[-8+16y-7y=1\]\[-8+9y=1\]\[-8+8+9y=1+8\]\[9y=9\]\[9y/9=9/9\]\[y=1\] Having found the value of \(y\), use your substitution equation \[x = -4 + 8y\]to find the value of \(x\) Finally, test your work by putting your values for \(x\) and \(y\) into both of your original equations and verifying that they produce true statements. You have to test all of the equations; a subset is not sufficient to demonstrate that your answer is correct, though it is sufficient to demonstrate that your answer is not correct :-)

OpenStudy (anonymous):

okay thanks!

OpenStudy (whpalmer4):

so, what do you get for x?